OpenStudy (anonymous):

solve for x: 16^x+4 = 3^2x-10...? help

OpenStudy (watchmath):

I suspect on the right should be 32^x

OpenStudy (anonymous):

oh, yes

OpenStudy (anonymous):

its 16^x+4 = 32^2x-10

OpenStudy (anonymous):

ughh cmon

OpenStudy (watchmath):

this is tough. Are you sure that this needs to be solved algebraically?

OpenStudy (anonymous):

yeah, it says solve for x on my paper

OpenStudy (watchmath):

well, but you can solve by graphing too...

OpenStudy (anonymous):

\[16^{x} + 4 = 32^{2x}-10\] or \[16^{x+4} = 32^{2x-10}\] or \[x ^{16} + 4 = 2x ^{32} -10\] which one is it?

OpenStudy (anonymous):

hi watchmath, im in a disagreement with amistre about integration region of revolution

OpenStudy (anonymous):

tje secound one audia

OpenStudy (anonymous):

the secound

OpenStudy (watchmath):

oh the second! then it is solvable I guess ....

OpenStudy (anonymous):

i want to revolve the region bounded by y = x+2 and y = x^2 about the y axis , this is not possible. since it annulls itself . amistre thinks it is possible

OpenStudy (anonymous):

you cant revolve a curve that is in Q2 or Q3 about the y axis, since then volume is not meaningful . the shell method for instance will output a negative ( i believe)

OpenStudy (anonymous):

forget it.. ill just try and figure this out.

OpenStudy (watchmath):

If it is the second equation, then we have \(4^{2x+8}=4^{5x-25}\) Hence \(2x+8=5x-25\) \(3x=33\) \(x=11\)

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

how did you turn it into 4^2x+8 = 4^5x-25?

OpenStudy (watchmath):

\(16^{x+4}=(4^2)^{x+4}=2^{2(x+4)}=4^{2x+8}\) Try to figure out the second one :D.